0.2 mol of HCL and 0.1 mol of barium chloride dissolved in water to produce a 500ml sol. The molarity of the CL- is?
1
Expert's answer
2017-06-05T08:40:09-0400
[〖Cl〗^- ]=[HCl]+2[Ba〖Cl〗_2 ]=0.2/0.5+2∙0.1/0.5=0.8 (M)